# A Comprehensive Survey on Deep Multimodal Learning with Missing Modality

### v2

@@ Abstract.
During multimodal model training and reasoning, data samples may miss certain modalities and lead to compromised model performance due to sensor limitations, cost constraints, privacy concerns, data loss, and temporal and spatial factors. This survey provides an overview of recent progress in Multimodal Learning with Missing Modality (MLMM), focusing on deep learning techniques. It is the first comprehensive survey that covers the historical background and the distinction between MLMM and standard multimodal learning setups, followed by a detailed analysis of current MLMM methods, applications, and datasets, concluding with a discussion about challenges and potential future directions in the field.
@@ 1. Introduction
(a) Multimodal Model Description
@@
Figure 1. (a) AI models can input and process data in a variety of different modalities; (b) Description of a three-modality scenario with full- and missing-modality samples. We abbreviate all “Modality” as “Mod” in all figures of this paper and visualize missing modalities with nearly transparent colors and dashed boxes.
@@
Multimodal learning has become a crucial field in Artificial Intelligence (AI). It focuses on integrating and analyzing various data types, including visual, textual, auditory, and sensory information (Figure 1(a)). This approach mirrors the human capacity to combine multiple senses for better understanding and interaction with the environment. Modern multimodal models leverage the robust generalization capabilities of deep learning to uncover complex patterns and relationships that uni-modality systems might not detect. This capability is advancing work across multiple domains, including computer vision. Recent surveys in these fields highlight the significant impact of multimodal approaches, demonstrating their ability to enhance performance and enable more sophisticated AI applications (Baltrušaitis et al., 2018; Zhao et al., 2024).
@@
However, multimodal systems often face the challenge of missing or incomplete data in real-world applications. This occurs due to various factors such as sensor malfunctions, hardware limitations, privacy concerns, environmental interference, and data transmission issues. As illustrated in  Figure 1(b), in a three-modality scenario, data samples can be categorized as either full-modality (containing information from all three modalities) or missing-modality (lacking data from one or more modalities entirely). These problems can arise at any stage from data collection to deployment, significantly impacting model performance. Real-world examples of this problem are prevalent across various domains. In affective computing, researchers (Cohen et al., 2004; Sebe et al., 2005) have encountered samples with only usable images or audio due to camera obstructions or excessive microphone noise. Similarly, in space exploration, NASA’s Ingenuity Mars helicopter (Donaldson, 2024) faced a missing modality challenge when its inclinometer failed due to extreme temperature cycles on Mars. To address this issue, NASA applied a software patch that modified the initialization of navigation algorithms (Team, 2022). In the field of medical AI, privacy concerns often result in certain modalities being unavailable for some data samples, creating inherent missing modalities in multimodal datasets (Zhang et al., 2023c). The unpredictability of real-world scenarios and diversity of data sources further compound this challenge. Consequently, developing robust multimodal systems that can perform effectively with missing modalities has become a crucial focus in the field.
@@
In this survey, we refer to the challenge of handling missing modalities as the ”missing modality problem.” We name solutions to this problem Multimodal Learning with Missing Modality (MLMM). This approach contrasts with the traditional setup that utilizes full modalities, which we name Multimodal Learning with Full Modality (MLFM). Specifically, in MLFM tasks, given a $H$-modality dataset, we are typically required to train a model can process and fuse information from all $H$ modalities to make predictions. During training and testing, samples with complete information from all $H$ modalities are used. In contrast, MLMM tasks use fewer than $H$ modalities during training or testing due to factors such as data collection limitations or constraints in deployment environments. The primary challenge in MLMM lies in dynamically and robustly handling and fusing information from any number of available modalities during training and testing, while maintaining performance comparable to that achieved with full-modality samples.
Figure 2. Our Methodology Taxonomy of Deep Multimodal Learning with Missing Modality Methods from four dimensions: modality augmentation, feature space engineering, architecture engineering, and model selection. “MLLMs” means multimodal large language models.
This survey encompasses recent advances in MLMM and its applications across various domains, including information retrieval, remote sensing, and robotic vision. We provide a fine-grained taxonomy of MLMM methodologies, application scenarios, and corresponding datasets. Our work expands upon existing MLMM surveys that focus on specific domains such as medical diagnosis (Zhou et al., 2023; Shah et al., 2023; Azad et al., 2022), sentiment analysis (Wagner et al., 2011), and multi-view clustering (Chao et al., 2021). By offering a comprehensive overview of current research and identifying promising directions for future work, this survey aims to contribute to the development of more robust and adaptable multimodal learning systems. These advancements are crucial for deploying intelligent systems in challenging environments, ranging from the harsh conditions of planetary exploration to the dynamic and unpredictable settings of everyday life. Main contributions of this survey are three-fold:
  1. (1)
    A comprehensive survey of MLMM applications across diverse domains, accompanied by an extensive compilation of relevant datasets, highlighting the versatility of MLMM in addressing real-world challenges.
  2. (2)
    A novel, fine-grained taxonomy of MLMM methodologies, with a multi-faceted categorization framework based on recovery strategies, integration stages, parameter-efficiency approaches, and attention mechanisms.
  3. (3)
    An in-depth analysis of current MLMM approaches, their challenges, and future research directions, contextualized within the proposed taxonomical framework.
@@
Paper Collection: In our literature search methodology, we primarily source papers from Google Scholar and major conferences/journals in Artificial Intelligence, Machine Learning, Computer Vision, Natural Language Processing, Audio Signal Processing, Data Mining, Multimedia, Medical Imaging, and Remote Sensing etc. We have compiled a total of 315 significant papers from the period spanning 2012 to 2024 September. Our search strategy involves using keywords such as “incomplete”, “missing”, “partial”, “absent”, and “imperfect”, coupled with terms like “multimodal learning”, “deep learning”, “representation learning”, and “neural network”.
@@
Survey Organization: Firstly, we explain the rough history of deep MLMM, background and motivation of this survey in Section 1. In Section 2, we categorize the methods of existing deep MLMM solutions from a methodological perspective, detailing them in four dimensions (Figure 2). We then summarize the current application scenarios and corresponding datasets used in this field in Section 3. In Section 4, we discuss the unresolved challenges and future directions of this field. Finally, in Section 5, we present our conclusions drawn from the exploration of deep MLMM.
2. Methodology
We classify and discuss existing deep MLMM methods based on our methodology taxonomy from four major dimensions: modality augmentation, feature space engineering, architecture engineering, and model selection.
2.1. Modality Augmentation
@@
We categorize modality augmentation methods addressing the missing modality problem at the original data of modality level into two types. The first is the modality composition method, which uses zero/random values, directly copied data from similar instances, or matching samples obtained via retrieval algorithms, composited with the missing-modality sample to form a full-modality sample. The second is the modality generation method, which generates the original data of the missing modality using generative models such as Auto Encoders (AEs) (Hinton and Zemel, 1993), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), or Diffusion Models (Ho et al., 2020).
@@
Figure 3. Zero/Random values composition method. If we assume modality 2 is missing, then this modality will be replace with zero/random values. “DNN” in all figures of this paper means different kinds of Deep Neural Networks.
@@
Figure 4. General idea of retrieval-based modality composition methods. We retrieval one or more samples by randomly selecting or using simple retrieval algorithms like KNN and its variants from past or same-category samples that have required missing modalities, and then compose them with the input missing-modality sample to form a full-modality sample.
2.1.1. Modality Composition Method
Removing missing-modality samples is common and simple as a multimodal dataset preprocessing method when the dataset has few missing-modality samples. But this method will reduce the dataset size and make some rare samples disappear when the dataset contains many missing-modality samples. The composition method is widely employed for its simplicity and effectiveness and also keep the dataset size. One typical method is the zero/random values composition method which replaces the missing modality data with zero/random values, as shown in Figure 3. In recent research (Chen et al., 2020; Sun et al., 2024a; Liu et al., 2023c; Malitesta et al., 2024), they are often used as baselines for comparison with proposed methods. About the missing sequential data problem, such as missing frames in videos, the similar Frame-Zero method (Parthasarathy and Sundaram, 2020) was proposed to replace the missing frames and composed with available frames. These methods are prevalent in typical multimodal learning training procedures. By using this, multimodal models can balance and integrate information from different modalities when making predictions. This helps prevent the model from relying too heavily on one or a few modalities to some extend, thereby enhancing the robustness. Additionally, these methods fail to generalize well when most samples in the dataset are the missing-modality samples. While, other composition method is based on retrieval algorithms (Figure 3), which involves filling in the original data of the missing modality by copying/averaging the original data from retrieved samples that have same modality and category. Some convenient methods randomly select a sample that has same category and required missing modality from other samples. The selected sample is then composed with the input missing-modality to form a full-modality sample for training. For example, researchers (Yang et al., 2024) proposed Modal-mixup to complete the training datasets by randomly complementing same-category samples with missing modalities. However, such methods cannot solve the missing modality problem during the testing phase. Regarding missing frames in streaming data such as videos, researchers proposed using Frame-Repeat (Parthasarathy and Sundaram, 2020) to make up for the missing frames by using past frames. Some work (Yang et al., 2024; Campos et al., 2015; Gong et al., 2021) also has attempted to use K-Nearest Neighbors (KNN) or its variants that can measure the similarity between samples to retrieve the best-matched samples for composition. For these matched samples, we can select the sample with the highest score or obtain the average values of these samples to supplement the missing modality data. Their experiments have shown that KNN-based methods generally perform better than above methods such as random selection and can handle missing modality during testing. But straightforward simple clustering often suffers from high computational complexity, sensitivity to imbalanced data, and high memory overhead. Also, retrieval-based modality combination methods are not applicable to pixel-level tasks such as segmentation and are only suitable for simple tasks (e.g., classification) because they may lead to model confusion. In addition, above all methods can complete datasets with missing modalities, it reduces the diversity of the dataset. This is especially problematic with datasets having a high modality missing rate (means most samples are missing-modality samples), as it increases the risk of overfitting to certain categories with few full-modality samples.
(a) Individual Modality Generation Method
(b) Unified Modality Generation Method

### v3

# Deep Multimodal Learning with Missing Modality: A Survey

@@ Abstract.
During multimodal model training and testing, certain data modalities may be absent due to sensor limitations, cost constraints, privacy concerns, or data loss, negatively affecting performance. Multimodal learning techniques designed to handle missing modalities can mitigate this by ensuring model robustness even when some modalities are unavailable. This survey reviews recent progress in Multimodal Learning with Missing Modality (MLMM), focusing on deep learning methods. It provides the first comprehensive survey that covers the motivation and distinctions between MLMM and standard multimodal learning setups, followed by a detailed analysis of current methods, applications, and datasets, concluding with challenges and future directions.
@@ 1. Introduction
(a) Trends in Missing Modality Publications (2014-2024 October)
@@
Figure 1. (a) The trend of papers published on deep multimodal learning with missing modality in the past 10 years. The number of publications has increased over time and has received widespread attention from the community. (b) Description of a three-modality scenario with full- and missing-modality samples. We abbreviate “Modality” as “Mod” in all figures of this paper and use dashed boxes with fading colors to represent missing modalities/modules.
@@
Multimodal learning has become a crucial field in Artificial Intelligence (AI). It focuses on jointly analyzing various data modalities, including visual, textual, auditory, and sensory information. This approach mirrors the human capacity to combine multiple senses for better understanding and interaction with the environment. Modern multimodal models leverage the robust generalization capabilities of deep learning to uncover complex patterns and relationships that single-modal systems might not detect. This capability is advancing work across multiple domains (Bayoudh et al., 2022; Hu et al., 2023; Streli et al., 2023). Recent multimodal learning surveys highlight the significant impact of multimodal approaches, demonstrating their ability to enhance performance and enable more sophisticated AI applications (Baltrušaitis et al., 2018; Zhao et al., 2024).
@@
However, real-world multimodal systems often face the challenge of having to handle cases where certain data modalities may be missing or incomplete. This occurs due to various factors such as sensor malfunctions, hardware limitations, privacy concerns, environmental interference, and data transmission issues. In recent years, more and more researchers have shown great interest in this field, which can be demonstrated by the bar chart in Figure 1(a). As illustrated in  Figure 1(b), in a three-modality scenario, data samples can be categorized as either full-modality (containing information from all three modalities) or missing-modality (lacking data from one or more modalities). These problems can arise at any stage from data collection to deployment, significantly impacting model performance. For instance, early affective computing researchers (Cohen et al., 2004; Sebe et al., 2005) found that when collecting facial and audio data, some image or audio samples were not useful due to excessive microphone noise or camera obstructions. This has forced them to propose audio-visual models that can handle the missing modality to recognize human emotional states. In the field of medical AI, privacy concerns and the challenges of obtaining certain data modalities during surgeries or invasive treatments often lead to missing modalities in multimodal datasets (Zhang et al., 2023c). Similarly, in space exploration, NASA’s Ingenuity Mars helicopter (Donaldson, 2024) faced a missing modality challenge when its inclinometer failed due to extreme temperature cycles on Mars. To address this issue, NASA applied a software patch that modified the initialization of navigation algorithms (Team, 2022). Moreover, the inherent structural or quantitative heterogeneity of sensors across different versions or brands of equipment can lead to variations in available modalities. This requires models to be capable of handling inputs with differing types of modalities in real-world applications. The unpredictability of real-world scenarios and diversity of data sources further compound this challenge. Consequently, developing robust multimodal systems that can perform effectively with missing modalities has become a critical focus in the field.
@@
Removing missing-modality samples is a common and simple multimodal dataset pre-processing strategy when the dataset has some missing-modality samples. But this method will waste the valuable information contained in missing-modality samples and cannot help models handle missing modality samples during testing, as it can only be used as a “temporary” solution for training. This has prompted many researchers to propose many multimodal learning methods for missing-modality smap. In this survey, we refer to the challenge of handling missing modalities for multimodal learning as the “missing modality problem.” We name solutions to this problem as Multimodal Learning with Missing Modality (MLMM). Those approaches contrast with the conventional setup that utilizes the full set of modalities, which we name as Multimodal Learning with Full Modality (MLFM). Specifically, in MLFM tasks, given an $\mathbf{N}$-modality dataset, we are typically required to train and test a model that can process and fuse all $\mathbf{N}$ modalities to make decisions. In contrast, MLMM tasks may use fewer than the $\mathbf{N}$ modalities available during training or testing due to factors such as data collection limitations or constraints in deployment environments. The primary challenge in MLMM lies in dynamically and robustly handling and fusing information from any number of available modalities during training and testing, yet maintaining performance comparable to that achieved with full-modality samples.
Figure 2. Our taxonomy of deep multimodal learning with missing modality methods. We categorize existing methods into two aspects: data processing and strategy design. Data Processing: we differentiate between modality imputation (handling at the modality data level) and representation-focused models (dealing with at the data representation level). Strategy Design: we distinguish between architecture-focused models (model architecture adjustments) and model combinations (combining multiple models externally). “MLLMs”: multimodal large language models.
This survey encompasses recent advances in MLMM and its applications across various domains, including information retrieval (Malitesta et al., 2024), remote sensing (Wei et al., 2023), robotic vision (Gunasekar et al., 2020), medical diagnosis (Zhou et al., 2023; Shah et al., 2023; Azad et al., 2022), sentiment analysis (Wagner et al., 2011), and multi-view clustering (Chao et al., 2021). We also introduce a fine-grained taxonomy of MLMM methodologies, application scenarios, and corresponding datasets. The main contributions of this survey are: (1) A comprehensive survey of MLMM methods across diverse domains, accompanied by an extensive compilation of relevant datasets, highlighting the versatility of MLMM in addressing real-world challenges. (2) A novel, fine-grained taxonomy of MLMM methodologies, with a multi-faceted categorization framework based on multi-modal integration stages and missing modality recovery strategies. (3) An in-depth analysis of current MLMM approaches, their challenges, and future research directions, contextualized within the proposed taxonomical framework.
@@
Paper Collection: In our literature search methodology, we primarily source papers from Google Scholar and major conferences/journals in AI, Machine Learning, Computer Vision, Natural Language Processing, Audio Signal Processing, Data Mining, Multimedia, Medical Imaging, and Remote Sensing. The collected papers are from, but not limited to, top-tier conferences (e.g., AAAI, IJCAI, NeurIPS, ICLR, ICML, CVPR, ICCV, ECCV, ACL, EMNLP, KDD, ACM MM, MICCAI, ICASSP) and journals (e.g., TPAMI, TIP, TMI, TMM, JMLR). Please refer to Section A in the Supplementary Materials for the full names of these conferences and journals. We have compiled a total of 315 significant papers from the period spanning 2012 to October 2024. Our search strategy involved using keywords such as “incomplete,” “missing,” “partial,” “absent,” and “imperfect,” combined with terms like “multimodal learning,” “deep learning,” “representation learning,” “multi-view learning,” and “neural networks.”
@@
Survey Organization: Firstly, we explain the background and motivation of this survey in Section 1. In Section 2, we introduce our taxonomy and categorize existing deep MLMM methods from a methodological perspective, detailing them in two aspects and four types (Figure 2). In the following Section 3 and Section 4, we introduce various methods from the aspects of model data processing and strategy design respectively. We then summarize current application scenarios and corresponding datasets used in this field in Section 6. In Section 7, we discuss unresolved challenges and future directions. Finally, we present our conclusions drawn from the exploration of deep MLMM in Section 8.
2. Methodology Taxonomy: An Overview
We review current deep MLMM methods from two key aspects: data processing and strategy design.
2.1. Data Processing Aspect
Methods that focus on exploring the data processing aspect of the model can be divided into Modality Imputation and Representation-Focused Models, according to whether the processing of missing modality occurs at the modality data level or the data representation level.
(1) Modality Imputation operates at the modality data level and fills in the missing information by compositing (Chen et al., 2020; Sun et al., 2024a; Parthasarathy and Sundaram, 2020; Yang et al., 2024, 2024; Campos et al., 2015) (modality composition methods) or generating (Chen et al., 2024b; Ma et al., 2021b; Arya and Saha, 2021; Wang et al., 2024a; Tran et al., 2017; Zhang et al., 2023b) absent modalities (modality generation methods) from available modalities. Those approaches are rooted in the idea that if a missing modality can be accurately imputed, downstream tasks can continue as if “full” modalities were available.
(2) Representation-Focused Models are designed to address missing modalities at the representation level. In some cases, the coordinated representation methods (Liang et al., 2019; Wang et al., 2020a; Ma et al., 2021c) impose some specific constraints on the representations of different modalities to help align the representations of different modalities in the semantic space, so that the model can be effectively trained even when facing missing modalities. Other representation-focused methods either generate the missing-modality representation using available data (Zheng et al., 2021; Li et al., 2024b; Hoffman et al., 2016; Ma et al., 2021b; Park et al., 2024) or combine the representations of existing modalities (Zhi et al., 2024; Sun et al., 2024a; Zhou et al., 2021; Delgado-Escano et al., 2021; Zheng et al., 2024; Zhang et al., 2024) to fill in the gaps.
2.2. Strategy Design Aspect
Methods that explore the strategy design aspect are based on models that can dynamically adapt to different missing-modality cases during training and testing through flexible adjustments of the model architecture (internal model architecture adjustment) and the combination of multiple models (external model combinations). We name them Architecture-Focused Models and Model Combinations.
(1) Architecture-Focused Models address missing modalities by designing flexible model architectures that can adapt to varying numbers of available modalities during training or inference. A key technique here is based on attention mechanisms (Ge et al., 2023; Yao et al., 2024; Gong et al., 2023; Mordacq et al., 2024; Radosavovic et al., 2024; Jang et al., 2024; Qiu et al., 2023), which dynamically adjusts the modality fusion and processing, allowing the model to handle any number of input modalities. Another approach is based on knowledge distillation (Wang et al., 2020b; Saha et al., 2024; Zhang et al., 2023c; Chen et al., 2024a; Wang et al., 2023a; Shi et al., 2024; Wang et al., 2023d), where the model is trained to accommodate missing modalities by transferring knowledge from full-modality models to those operating with incomplete data or distilling between different branches inside the model. Additionally, graph learning-based methods (Zhao et al., 2022; Zhang et al., 2022; Lian et al., 2023; Malitesta et al., 2024) exploit the natural relationships between modalities, using graphs to dynamically fuse and process available modalities while compensating for missing ones. Finally, MLLMs (Zhan et al., 2024; Wu et al., 2023a; Li et al., 2023c) also play a crucial role in this category, as their ability to handle long contexts and act as feature processors enables them to accept and process representations from any number of modalities. These architectural strategies collectively allow models to maintain performance even when dealing with incomplete multimodal inputs.
(2) Model Combinations tackle missing modality problems by employing strategies that leverage multiple models or specialized training techniques. One approach is to use dedicated training strategies (Zeng et al., 2022; Chen et al., 2022; Xue and Marculescu, 2023) tailored for different modality cases, ensuring that each case is trained for optimal performance. Another approach involves ensemble methods (Hu et al., 2020; Wang et al., 2021; Lee et al., 2023b), where models trained on either partial/full sets of modalities are combined, allowing the system to select the most suitable model based on the available modalities to do joint predictions. Additionally, discrete scheduler methods (Wu et al., 2023b; Shen et al., 2024; Surís et al., 2023) can incorporate various downstream modules to flexibly process any number of modalities and handle specific tasks. These schedulers intelligently select and combine the outputs of multiple models or modules to manage missing-modality scenarios, offering a versatile solution for multimodal tasks.
Our taxonomy (Figure 2) can reflect different aspects and levels of multimodal learning—ranging from modality data to data representations, architectural design, and model combinations—each providing a distinct way to approach the problem of missing modalities based on the task requirements and available resources.
3. Methodologies in Data Processing Aspect
3.1. Modality Imputation
@@
Modality imputation refers to a technique used by MLMM methods that fills in missing-modality samples or generates missing modalities to complete the dataset with missing modalities by performing various transformations or operations on existing modalities. Modality imputation methods addressing the missing modality problem at the data modality level can be categorized into two types. (1) Modality composition methods use zero/random values or data copied from similar instances as the input for the missing modality data. The data produced via these methods to represent the missing data are then composited with the data from the available modalities to form a “full”-modality sample. (2) Modality generation methods generate the missing modality data using generative models such as Auto-Encoders (Hinton and Zemel, 1993), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), or Diffusion Models (Ho et al., 2020). The generated data is then composited with the data from the available modalities to form a “full”-modality sample. We provide more details about these two methods in the next sub-sections.
@@
Figure 3. Zero/Random values composition methods. If we assume modality 2 is missing, then this modality will be replaced with zero/random values. “DNN” in all figures of this survey means different kinds of deep neural networks.
@@
Figure 4. Retrieval-based modality composition methods search for one or more samples by randomly selecting or using simple retrieval algorithms like KNN, or its variants, from same-category samples that have the required missing modalities, and then compose them with the input missing-modality sample to form a “full”-modality sample.
3.1.1. Modality Composition Methods
Modality composition methods are widely employed for its simplicity and effectiveness to maintain the original dataset size. Zero/Random Values Composition Methods represent a type of modality composition method that replaces a missing modality with zeros or random values, as shown in Figure 3. In recent research (Chen et al., 2020; Sun et al., 2024a; Liu et al., 2023c; Malitesta et al., 2024), they are often used as a baseline method for comparison with other more sophisticated methods. For the missing sequential data problem, such as missing frames in videos, the similar Frame-Zero method (Parthasarathy and Sundaram, 2020) was proposed to replace the missing frames. These methods are prevalent in typical multimodal learning procedures and can be used to balance and integrate information from different modalities when making predictions. This prevents the model from over-relying on dominant modalities available for each sample, enhancing its robustness by encouraging a more balanced integration of information across all available modalities.
Retrieval-Based Representation Composition Methods (Figure 4) represent another type of modality composition method, which replaces the missing modality data by copying or averaging data from retrieved samples with the same classification. Some other methods randomly select a sample that has the same classification and required missing modality from other samples. The selected modality data is then composed with the missing-modality sample to form a full-modality sample for training. But those retrieval-based modality composition methods are not applicable to pixel-level tasks, such as segmentation, and are only suitable for simple tasks (e.g., classification) because they may lead to the overfitting of noisy data, if mismatched samples are combined. For example, Yang et al. (Yang et al., 2024) proposed Modal-mixup to complete the training datasets by randomly complementing same-category samples with missing modalities. However, such methods cannot solve the missing modality problem during the testing phase because they rely on known labels of training data samples. In some multimodal streaming data classification tasks, such as audio-visual expression recognition, video streams may experience frame drops due to network communication packet loss, etc. Frame-Repeat (Parthasarathy and Sundaram, 2020) was proposed to to make up for the missing frames by using past frames.
Other methods (Yang et al., 2024; Campos et al., 2015) also used K-Nearest Neighbors (KNN), or its variants, to retrieve the best-matched samples for composition. For these matched samples, they select the sample with the highest score or obtain the average values of these samples to supplement the missing modality data. Their experiments have shown that KNN-based methods generally perform better than the above methods, and can handle missing modality during testing. Nevertheless, such KNN retrieval-based modality composition methods often suffer from high computational complexity, sensitivity to imbalanced data, and significant memory overhead.
All methods above can complete datasets with missing modalities, but they reduce the diversity of the dataset because they may introduce duplicated training samples. This is especially problematic with datasets having a high rate of modality missing, where most samples have missing-modality data, as it increases the risk of overfitting to certain classes with few full-modality samples, if they pad missing-modality samples with duplicated ones.
(a) Individual Modality Generation Methods
(b) Unified Modality Generation Methods